OpenStudy (anonymous):

sin^4 θ − cos^4 θ = sin^2 θ − cos^2 θ asked to verify the identity

jimthompson5910 (jim_thompson5910):

Hint: x^4 - y^4 = (x^2)^2 - (y^2)^2 x^4 - y^4 = (x^2-y^2)(x^2+y^2)

jimthompson5910 (jim_thompson5910):

not sure why he deleted all that, but he's right

OpenStudy (anonymous):

haha, yea, me either.

OpenStudy (anonymous):

so I say sin^4-cox^4=(sin^2)^2(cox^2)^2 and =(sin^2-cos^2)(sin^2+cos^2) ?

jimthompson5910 (jim_thompson5910):

yes, then use the idea that sin^2+cos^2 = 1

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

I'm assuming you meant to say sin^4-cox^4=(sin^2)^2 - (cox^2)^2

OpenStudy (anonymous):

oh, yea

OpenStudy (anonymous):

hmm, I'm not sure I see how to break that down to just sin^2 and cos^2

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (e.mccormick):

This type of thing is all about starting out with a good list of trig identities and seeing how few times you need to use it. The more times you do identities, the less you use the references. Eventually you just know them from using them so many times.

OpenStudy (anonymous):

yea, I've been trying to memorize the learn the formulas well

OpenStudy (e.mccormick):

I am thinking power reduction and/or sum and difference would be good ones to look at. Don't have the whole thing in my head yet, but from what I remember, I see those as a good start.

OpenStudy (anonymous):

ok, Ill look into that also, thank you!

OpenStudy (e.mccormick):

Now for sleep. I work in the morning... and have class in the afternoon. Whee! I hope what Jim and I pointed to leads somewhere good.

OpenStudy (anonymous):

ohk lets start by solving the left hand side that is sin^4 d-cos^4 d (here 'd' is an angle just like theta tht u hv assumed in the ques) alrt u must be knowing the identity a^2-b^2=(a+b)(a-b) so sin^4 d -cos^4 d=(sin^2 d)^2 -(cos^2 d)^2 =[(sin^2 d)+(cos^2 d)][(sin^2 d)-(cos^2 d)] since [(sin^2 d)+(cos^2 d)]=1 so ans is [(sin^2 d)-(cos^2 d)] which is equal to the RHS :p hence proved :)